OpenStudy (anonymous):

The 5th term of an arithmetic sequence described by the rule is tn=19-3n ? Explanation?

OpenStudy (anonymous):

substitute n=5 in the formula above, what do you get?

OpenStudy (anonymous):

5=19-3n? Is it reverse bodmass because you are solving for n?

OpenStudy (anonymous):

term 5 = 19 - 3*5 = 4 thats your answer

OpenStudy (anonymous):

Ya im not sure how to solve it.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Welcome do you know what an arithmetic sequence is?

OpenStudy (anonymous):

So you don't substitute 5 in tn?

OpenStudy (anonymous):

tn stands for Term number n , you can say this \[t_{5} = 19 - 3*5 = 4\]

OpenStudy (anonymous):

@fliptricks1 are you ok with the above?